# RUBIES: JWST/NIRSpec Confirmation of an Infrared-luminous, Broad-line Little Red Dot with an Ionized Outflow

### v1

@@ 4.5 Mass Outflow Rate
There is a strong positive correlation between the radius and the mass outflow rate. The lowest density models reach radii of pc with corresponding mass outflow rates of . On the other hand, the highest density models imply radii of 0.1 pc, which would place the absorber within the AGN broad line region, with corresponding mass outflow rates 0.01 M yr. To constrain this better would require another estimate of the absorber size, more absorption lines, or both.
@@ 5 Discussion and Conclusions
We present spectroscopic observations of a bright IR-luminous broad-line LRD using NIRSpec/Prism and G395M onboard JWST. We perform spectrophotometric modeling to infer the properties of the host galaxy and the AGN, and decompose the emission and absorption complexes to constrain the gas kinematics and measure the black hole mass. Additionally, we take a first look at the mass outflow rate as implied from the He i 1.083m absorption. Below we discuss the key findings in this work.
@@ 5.1 A Spectroscopically Confirmed Broad-Line AGN with a Deficit of Hot Dust
We have developed a novel method to model the host galaxy and AGN within the Prospector Bayesian inference framework (Johnson et al., 2021). Our preferred model is a composite AGN+host model, in which the UV and the Balmer break are explained by galaxy light, and the AGN continuum from the outer accretion disk dominates in the optical and mid-IR. We find that explaining the rest-frame optical continuum light with pure galaxy continuum is possible, but yields a very high mass of . While unresolved in the rest-frame optical, its measured half light radius ( kpc at F277W) would yield an extreme stellar surface density kpc; even a conservative upper limit of one pixel on its size would yield a very high stellar density kpc (in some tension with theoretical density limits from stellar feedback; Hopkins et al., 2010; Grudić et al., 2019) and is inconsistent with the dynamical mass inferred from the narrow Pa- and Pa- emission lines. Further, a galaxy interpretation also would lead to a very high H rest-frame EW of Å, which is inconsistent with the inferred older stellar populations and also has not been seen in known AGNs (Vanden Berk et al., 2001; Stern & Laor, 2012). Likewise, fitting a pure AGN model fails to reproduce the spectral shape around the Balmer break region (see Appendix A).
@@
The flat (in ) spectral shape in the mid-IR also presents a significant challenge to our composite AGN and host galaxy picture since a near-ubiquitous component of AGN at all redshifts is a hot dusty torus that begins to dominate AGN emission longward of 2 . That component is not observed in this source. In fact, a much larger sample of sources detected in the JADES fields tells the same basic story; while the rest-frame optical appears to require significant obscuration, there is no evidence of emission from the torus itself out to in the rest frame (Williams et al., 2023; Pérez-González et al., 2024). In the majority of cases, with the exception of one object from Matthee et al. (2023), this larger sample do not have measurements of broad H, and so RUBIES-BLAGN-1 provides a special case where we can strongly argue for an AGN component despite the lack of hot dust, perhaps refuting the assumption made in Pérez-González et al. (2024) that these sources are best explained by dusty star formation in the majority of cases. However, RUBIES-BLAGN-1 provides clear confirmation that it is possible to both have an unambiguous broad-line AGN dominate the red continuum in the rest frame optical and simultaneously lack the rising hot dust continuum in the mid-infrared.
@@ 5.4 Final Remarks
The red objects discovered by JWST continue to challenge the standard views surrounding the evolution of black holes and galaxies. This paper presents an in-depth analysis of an LRD that has both broad lines and JWST/MIRI detections, an occurrence that was observed only once previously. While it remains uncertain whether RUBIES-BLAGN-1 represents the typical properties of LRDs in terms of the inferred characteristics of the host galaxy, AGN, and outflows, this work demonstrates a promising prospect for understanding the nature of the red sources. The RUBIES program will obtain a statistical sample in the near future, and we plan to extend our modeling framework to perform a systematic investigation, which will contribute to complete an important chapter of the narrative of black hole growth.
@@ Acknowledgments
B.W. and J.L. acknowledge support from JWST-GO-04233.009-A. R.L.D. is supported by the Australian Research Council through the Discovery Early Career Researcher Award (DECRA) Fellowship DE240100136 funded by the Australian Government. The Cosmic Dawn Center is funded by the Danish National Research Foundation (DNRF) under grant #140. This research was supported by the International Space Science Institute (ISSI) in Bern, through ISSI International Team project #562 (First Light at Cosmic Dawn: Exploiting the James Webb Space Telescope Revolution). Computations for this research were performed on the Pennsylvania State University’s Institute for Computational and Data Sciences’ Roar supercomputer. This publication made use of the NASA Astrophysical Data System for bibliographic information.
@@
A Alternative SED Models
@@ A.2 Dust
A.3 Fitting the Rest Optical Spectrum Only

### v2

@@ 4.6 Mass Outflow Rate
There is a strong positive correlation between the radius and the mass outflow rate. The lowest density models reach radii of pc with corresponding mass outflow rates of . On the other hand, the highest density models imply radii of 0.1 pc, which would place the absorber within the AGN broad line region, with corresponding mass outflow rates 0.01 M⊙ yr-1. To constrain this better would require another estimate of the absorber size, more absorption lines, or both.
@@ 5 Discussion and Conclusions
In this paper, we present spectroscopic observations of a bright IR-luminous broad-line LRD using NIRSpec/Prism and G395M onboard JWST. We perform spectrophotometric modeling to infer the properties of the host galaxy and the AGN, and decompose the emission and absorption complexes to constrain the gas kinematics and measure the black hole mass. Additionally, we take a first look at the mass outflow rate as implied from the He i 1.083m absorption. Below we discuss the key findings in this work.
@@ 5.1 A Spectroscopically Confirmed Broad-line AGN with a Deficit of Hot Dust
We have developed a novel method to model the host galaxy and AGN within the Prospector Bayesian inference framework (Johnson et al., 2021). Our preferred model is a composite AGN and host model, in which the UV and the Balmer break are explained by galaxy light, and the AGN continuum from the outer accretion disk dominates in the optical and mid-IR. These results are further corroborated by our morphological analysis of RUBIES-BLAGN-1, in which we find the source to be marginally resolved in the rest-frame UV, but unresolved at rest-frame optical wavelengths (Appendix A).
Explaining the rest-frame optical continuum light with pure galaxy continuum is possible, but yields a very high mass of . While unresolved in the rest-frame optical, its measured half light radius ( kpc at F277W) would yield an extreme stellar surface density kpc-2; even a conservative upper limit of one pixel on its size would yield a very high stellar density kpc-2 (in some tension with theoretical density limits from stellar feedback; Hopkins et al., 2010; Grudić et al., 2019) and is inconsistent with the dynamical mass inferred from the narrow Pa- and Pa- emission lines. Further, a galaxy interpretation also would lead to a very high H rest-frame EW of Å, which is inconsistent with the inferred older stellar populations and also has not been seen in known AGNs (Vanden Berk et al., 2001; Stern & Laor, 2012). Likewise, fitting a pure AGN model fails to reproduce the spectral shape around the Balmer break region (Appendix B).
@@
The flat (in ) spectral shape in the mid-IR also presents a significant challenge to our composite AGN and host galaxy picture since a near-ubiquitous component of AGN at all redshifts is a hot dusty torus that begins to dominate AGN emission longward of 2 . That component is not observed in this source. In fact, a much larger sample of sources detected in the JADES fields tells the same basic story; while the rest-frame optical appears to require significant obscuration, there is no evidence of emission from the torus itself out to in the rest frame (Williams et al., 2023; Pérez-González et al., 2024). In the majority of cases, with the exception of one object from Matthee et al. (2023), this larger sample do not have measurements of broad H, and so RUBIES-BLAGN-1 provides a special case where we can strongly argue for an AGN component despite the lack of hot dust, perhaps refuting the assumption made in Pérez-González et al. (2024) that these sources are best explained by dusty star formation in the majority of cases. However, RUBIES-BLAGN-1 provides clear confirmation that it is possible to both have an unambiguous broad-line AGN dominate the red continuum in the rest frame optical and simultaneously lack the rising hot dust continuum in the mid-IR.
@@ 5.4 Final Remarks
The red objects discovered by JWST continue to challenge the standard views surrounding the evolution of black holes and galaxies. This paper presents an in-depth analysis of an LRD that has both broad lines and JWST/MIRI detections, an occurrence that was observed only once previously. The lower redshift of RUBIES-BLAGN-1 provides a better coverage at the rest-frame mid-IR, making the lack of torus emission more evident. While it remains uncertain whether RUBIES-BLAGN-1 represents the typical properties of LRDs in terms of the inferred characteristics of the host galaxy, AGN, and outflows, this work demonstrates a promising prospect for understanding the nature of the red sources. The RUBIES program will obtain a statistical sample in the near future, and we plan to extend our modeling framework to perform a systematic investigation, which will contribute to an important chapter of the narrative of black hole growth.
@@ Acknowledgments
B.W. and J.L. acknowledge support from JWST-GO-04233.009-A. R.L.D. is supported by the Australian Research Council through the Discovery Early Career Researcher Award (DECRA) Fellowship DE240100136 funded by the Australian Government. T.B.M. was supported by a CIERA postdoctoral fellowship. The Cosmic Dawn Center is funded by the Danish National Research Foundation (DNRF) under grant #140. This research was supported by the International Space Science Institute (ISSI) in Bern, through ISSI International Team project #562 (First Light at Cosmic Dawn: Exploiting the James Webb Space Telescope Revolution). Computations for this research were performed on the Pennsylvania State University’s Institute for Computational and Data Sciences’ Roar supercomputer. This publication made use of the NASA Astrophysical Data System for bibliographic information.
@@
A Morphology
We perform a morphological analysis of RUBIES-BLAGN-1 in the short wavelength filter F115W and the long wavelength filter F277W, which probe the rest-frame UV (m) and rest-frame optical (m), respectively. We create a custom mosaic for the F115W imaging, largely following the procedure described in Section 2, but drizzled to an improved pixel scale of 0.02″. Empirical PSFs are constructed from non-saturated stars on the 0.02″ mosaic for F115W, and the 0.04″ mosaic for F277W (A. Weibel et al. in prep).
We first measure the radial surface brightness profiles using the photutils.profiles module (Bradley et al., 2023) with circular apertures that are one pixel wide, ranging from pixels. As seen from Figure 8, the surface brightness profile of RUBIES-BLAGN-1 extends beyond the profile of the PSF in the rest-frame UV (F115W), indicating that it is indeed spatially extended. Conversely, in F277W the brightness profile of RUBIES-BLAGN-1 and the PSF show no significant differences, indicating that it is a point source at longer wavelengths.
Second, we utilize pysersic (Pasha & Miller, 2023) to fit a Sérsic profile (Sersic, 1968) to the images. We assume a uniform prior for the effective radius (between 0.25 to 10 pixels) and a uniform prior on Sersic index (between 0.65 to 4). We use the No-U-Turn Sampler (NUTS; Hoffman et al. 2014; Phan et al. 2019) to explore the posterior distribution of the morphological parameters with four chains and one thousand warm-up steps, and then five thousand sampling steps for each filter.
Consistent with our findings based on the surface brightness profiles, we find RUBIES-BLAGN-1 to be marginally resolved with an effective radius in the F115W filter. The Sérsic index is largely unconstrained but the posterior displays a slight preference for . Focusing on the pixel region centered on RUBIES-BLAGN-1, the maximum a-posteriori Sérsic model shows a lower by 15% compared to the value when modeling it solely as a point source. In F277W, RUBIES-BLAGN-1 is unresolved with an inferred . The Sérsic model shows no improvement in the values compared to modelling it as a point source. Both Sérsic models and marginalized posteriors for the effective radii are included in Figure 8.
Furthermore, we draw 500 posterior samples from the F115W Sersic fit, convolve the model images with the F277W PSF, and measure the surface brightness profiles. The median and the widths are over-plotted in blue in Figure 8 (b). The resulting model profile is slightly extended compared to the PSF: therefore, if the rest-frame optical light of RUBIES-BLAGN-1 were as extended as the rest-frame UV, then it should also be marginally resolved in F277W. Crucially, we do not observe such a difference, which implies that the UV and rest-frame optical indeed trace different physical components.
To summarize, we find that RUBIES-BLAGN-1 is marginally resolved in F115W, but unresolved in F277W. This analysis is independent of our SED modeling, but, interestingly, consistent with our SED modeling results (§ 4.3): the emission in the rest-frame UV originates from the host galaxy and thus is extended spatially, whereas the rest-frame optical emission is dominated by the AGN that is point-like in nature.
B Alternative SED Models
@@ B.2 Dust
B.3 Fitting the Rest Optical Spectrum Only